# microtypeUnable to apply patch

### v1

@@ 4 Results
Table 1 compares the performance of the baseline, unaligned model to the supervised fine-tuned model (SFT) and to the SFT model further aligned with a preference optimization algorithm (DPO).
@@
We observed a clear improvement across the board in empathy and similarity metrics provided by the SFT step. However, the more intensely the model is fine-tuned, the more it over-fits on the ED dataset, and the general performance drops, as shown with language understanding metrics in Figure 1. Mitigating this, but not avoiding it, is possible through careful hyperparameter optimization. First, as Tunstall et al., 2023 suggests, we limited the training to one epoch. The key to reducing over-fitting was using a large lora_rank ($r$) and low lora_alpha ($\alpha$) relative to the rank. The large $\alpha$ supports retention of the original model’s abilities, and the rank $r$ controls the impact size of the adapter’s fine-tuned weights. Figure 2 shows the empathy metrics saturating with $\alpha\approx\frac{r}{4}$. Furthermore, we observed that increasing the rank above 1024 brings diminishing returns: the computational efficiency of LoRA vanishes while the over-fitting problem remains. We use the best configuration jointly for empathy and generalization ($r\!=\!1024,\alpha\!=\!256,lr\!=\!1e\!-\!5$) to report SFT results and to perform DPO experiments.
Figure 1: SFT: Impact of LoRA $\alpha$ on the MMLU score. Trained with: learning_rate=1-e5, batch_size=64.
Figure 2: SFT: Impact of LoRA adapter rank on the diff-Epitome score during SFT. Trained with: learning_rate=1-e5, batch_size=64.
@@
The preference alignment with DPO improved the empathy metrics over the SFT while retaining the model’s general performance. Nevertheless, its hyperparameters need to be set suitably to achieve this. Unsuitable hyperparameter configuration leads to training instability in addition to over-fitting (see Figure 3). Notably, we faced problems with the stability of the training introduced by the dataset construction process: random selection of rejected completions from the group of dialogues labeled with polar opposite emotions. We solved it by training for more epochs: re-drawing new randomly selected rejected completions for the same preferred ones for each epoch. The optimum configuration for high empathy scores while controlling over-fitting was training for three epochs, same as in Tunstall et al., 2023, with low $\beta$ ($0.025$) but a relatively high learning rate ($51e\!-\!$ 5). We did not optimize BERTscore in SFT or DPO training, and results indicate no correlation with the diff-Epitome metric.
@@
Figure 3: Stability of the DPO training measured by the spread of metric scores between multiple runs with the same hyperparameters. A smaller spread signifies higher training stability. We measure the mean and spread of the diff-Epitome metric, averaged across its three dimensions (ER, EX, IP: lower is better). We observe that setting $\beta\!>>\!0.05$ leads to over-fitting on the preferred completions from the training set.
@@ 5 Conclusion
We introduced a new method for constructing preference datasets to align large language models (LLMs) for empathetic response generation while maintaining their overall language understanding capabilities. We validate our findings using the EmpatheticDialogues dataset and evaluate using the diff-Epitome and BERTscore metrics for empathy and standard language understanding benchmarks. Our method achieves the desired alignment of LLMs and provides a robust foundation for further enhancements, such as through prompt engineering. Furthermore, our method is directly applicable to other datasets with emotion labels.
@@
Empathetic Responses
Lahnala et al., 2022 argue that most NLP research defines empathy loosely as understanding and appropriately responding to others’ emotions, focusing mainly on detecting sentiment, emotions, or supportive interactions in text as indicators of empathy. This perspective assumes that systems achieving emotional recognition and responding in line with the target’s sentiment are empathetic. However, this approach overlooks the critical aspect of cognitive empathy, which involves understanding another person’s perspective, a gap highlighted by established human empathy theories Debnath and Conlan (2023).
Instead, we use a fully data-driven approach. We utilize the EmpatheticDialogs dataset as a source of the ground-truth empathetic responses. Even though there is criticism (Debnath and Conlan, 2023), dialogues in this dataset are written by human conversation partners and thus align well with real-world empathetic interactions.
@@ Appendix A Training Details
A.1 Model Settings
We trained the models using the HuggingFace Transformers library. In the SFT step, we applied standard sequence-to-sequence training with cross-entropy loss on tokens from the user prompts (not on the tokens from "replies"). In both the SFT and DPO training steps, we optimized the model with AdamW optimizer and effective batch size of 64. We used learning rate of 1e-5 without warmup and a cosine lr decay. The models were trained in bf16 precision.
A.2 Hardware
To train our models, we used two NVIDIA A100 80GB GPUs. We trained LoRA adapters with ranks ranging from 16 to 2048 for models with 7BM parameters. The total training wall time, including preliminary experiments, was 17 days.
A.3 DPO Preference Dataset

### v2

@@ 4 Results
4.1 Automatic Evaluation
Tables 2 and 3 compare the baseline, unaligned model to the supervised fine-tuned model (SFT) and to the SFT model further aligned with a preference optimization algorithm (DPO) on the main empathy and language understanding metrics.
@@ 4.1 Automatic Evaluation
We observed a specialization effect provided by the SFT step, which improved two dimensions of diff-Epitome and BERTscore similarity. However, the more intensely the model is fine-tuned, the more it over-fits on the ED dataset, and the language understanding performance drops, as shown in Figure 1. It is possible to mitigate this but not avoid it through careful hyperparameter optimization. First, as Tunstall et al., 2023 suggests, we limited the training to one epoch. The key to reducing over-fitting was using a large lora_rank ($r$) and low lora_alpha ($\alpha$) relative to the rank. Rank controls the size of adapter matrices, and $\alpha$ scales the impact of the adapter weights on the base model. Large $\alpha$ supports retention of the original model’s abilities, and rank $r$ controls the impact size of the adapter’s fine-tuned weights. Figure 2 shows the empathy metrics saturating with $\alpha\approx\frac{r}{4}$. Furthermore, we observed that increasing the rank above 1024 brings diminishing returns: the computational efficiency of LoRA vanishes while the over-fitting problem remains. We use the best configuration jointly for empathy and generalization ($r\!=\!1024,\alpha\!=\!64,lr\!=\!1e\!-\!5$) to report SFT results and subsequently to perform DPO training.
Figure 1: Supervised fine-tuning with LoRA: impact of the hyperparameter $\alpha$ on the MMLU score and Open LLM Leaderboard accuracy. Trained with: lr=1-e5, batch_size=64, rank=1024.
Figure 2: Supervised fine-tuning with LoRA: impact of the hyperparameter $\alpha$ on the diff-Epitome score. Trained with: lr=1-e5, batch_size=64, rank=1024.
@@
The preference alignment with DPO improved the empathy metrics over the SFT while retaining the model’s general performance. Nevertheless, its hyperparameters need to be set suitably to achieve this. Unsuitable hyperparameter configuration leads to training instability in addition to over-fitting (see Figure 3). Notably, we faced problems with the stability of the training introduced by the dataset construction process: random selection of rejected completions from the group of dialogues labeled with polar opposite emotions. We solved it by lowering $\beta$ and learning rate and training for more epochs: re-drawing new randomly selected rejected completions for the same preferred ones for each epoch. The optimum configuration for high empathy scores while controlling over-fitting was training for three epochs, same as in Tunstall et al., 2023, with low $\beta$ ($0.05$) and moderate learning rate ($61e\!-\!$ 6).
@@
Figure 3: Stability of the DPO training measured by the spread of scores between multiple runs with the same hyperparameters, different seed and new completion draw; smaller spread signifying more stability. We measure the mean and spread of diff-Epitome, averaged across dimensions (ER, EX, IP): lower is better.
Figure 4: Evaluations based on Lee et al. (2024)’s work: specificity with mean NIDF See et al. (2019).
Refer to caption Refer to caption Refer to caption
Figure 5: Results of word choice metrics: intensity, valence, and arousal (IVA). Difference between the prompt and generated response score i.e., $\text{score}(\text{prompt})-\text{score}(\text{utterance})$. The metric is further detailed in Appendix C.
[…]
Human Response 2526 (99.4%) 2746 / 22279 (12.3%) 881 (34.7%) 65.64% 370
baseline: Zephyr-7B-sft-full 2529 (99.6%) 5339 / 29080 (18.4%) 788 (31.0%) 44.53% 31
length-controlled baseline 2326 (91.6%) 2625/13425 (19.6%) 787 (31.0%) 53.33% 35
Zephyr-7B + SFT 2390 (94.1%) 2085 / 9437 (22.1%) 741 (29.2%) 47.90% 48
Zephyr-7B + SFT + DPO 2538 (99.9%) 2927 / 20074 (14.6%) 806 (31.7%) 60.36% 256
Table 4: Diversity metrics based on response-tries from Lee et al. (2024). # Templates refers to the number of unique responses generated by each system. Compression ratio refers to the ratio between the size of tries made before and after the folding operations. A higher compression ration implies lower repetition. The metric is detailed in Appendix C.
Diversity, Sensitivity, and Emotional Valence
Across the diversity metrics shown in Table 4, the human scores outperform all models, except in the case of # Templates. DPO responses show the best results among the models for all other metrics. Generating more diverse responses is desirable because models tend to resort to generating generic responses like “I am so sorry to hear that,” which may bolster their performance on metrics that evaluate individual examples. Lee et al. (2023) shows that Epitome and other representation-based models can take these shortcuts, which goes against the context-dependent nature of empathy.
Table 5 shows specificity, intensity, valence, and arousal. We performed Welch’s t-tests to determine the significance of differences between each model and human values. For specificity, each model differs significantly from human responses, as shown in Figure 4. The Zephyr models have lower specificity ($t=-50.8$, $p<0.001$; $t=-2.1$, $p=0.039$) and baseline with higher ($t=2.6$, $p=0.012$). The models’ specificities differ significantly from each other ($p<0.001$); SFT has the lowest and the baseline the highest, though there is an ideal point beyond which specificity does not improve model performance Coll et al. (2017).
Specificity Word choice
[…]
Human Response $\mathbf{0.30}\pm 0.07$ $0.28$ $0.14$ $0.18$
baseline: Zephyr-7B $\mathbf{0.30}\pm 0.05$ $\mathbf{0.24}$ $\mathbf{0.11}$ $\mathbf{0.15}$
len-control baseline $0.26\pm 0.05$ $0.26$ $0.12$ $\mathbf{0.15}$
SFT $0.22\pm 0.04$ $0.29$ $0.13$ $0.18$
SFT + DPO $0.29\pm 0.06$ $0.27$ 0.130.13$\ddagger$0.13 ‡ 0.160.16$\dagger$0.16 †
Table 5: Evaluations based on Lee et al. (2024): specificity with mean NIDF and word choice (IVA). Values closer to 0 indicate the response had a closer IVA score to the prompt. * Indicates significant improvement over Human Response with $p<0.00001$, $\dagger$ with $p<0.001$, and $\ddagger$ with $p<0.01$ using permutation tests. The metric is detailed in Appendix C.
Figure 5 shows the intensity-valance-arousal (IVA) metrics results. Desirable IVA metrics are closer to zero, signifying that IVA levels of responses are more similar to the prompts. The DPO model and baseline have significantly lower valence and arousal than humans, and the baseline has significantly lower intensity ($p<0.05$). Between models, the baseline is significantly lower than the SFT and DPO models for all IVA metrics, and DPO is lower than SFT for arousal and intensity. Ultimately, the DPO model shows the most similar performance to human responses overall. For the absolute values of IVA metrics (as in Lee et al. (2024)) and complete distributions of the metrics, see Appendix C.
4.2 Human Evaluation
Figure 6 summarizes the comparative human evaluation. We evaluate samples from the EmpatheticDialogues corpus, and ask raters to evaluate empathetic engagement in four dimensions. Interestingly, samples directly taken from the corpus (human in Fig. 6) were rated lower than their generated counterparts, while the Zephyr baseline model was considered most empathetic.
All of the models and the corpus sample share similar average scores in Consistency and Fluency ($\approx 0.92$).
The DPO model was rated as equally or more empathetic than the length-controlled baseline and SFT models in most dimensions and consistently scored higher in response naturalness. However, the baseline outperformed the DPO and SFT models in situational appropriateness, likely due to a learned bias from fine-tuning by focusing on ’situational empathy’ in the ED corpus. However, the difference was small but statistically significant.
The ratings are relatively independent of the emotion label that they were sampled from, and dimension ratings are not correlated. In summary, across independent dimensions of empathy, human raters showed a mixed preference leaning toward the DPO model’s responses (excluding the Zephyr baseline, which was overwhelmingly preferred). Consistently low ratings across empathy dimensions for samples from the original corpus also echo concerns about corpus quality from Debnath and Conlan (2023).
Figure 6: Distribution of human ratings along four dimensions of empathetic response using continuous scales (0-100). See Appendix D for the dimension descriptions.
4.3 Response Length: Observations
The mean length of ground truth and the DPO responses is 65 and 63 characters respectively, 55 for SFT, and 152 for the Zephyr-7B baseline. The long-response bias affects each of the automated metrics and also affects the human evaluation, which we were unable to fully mitigate by the survey design. This is not surprising, as there is evidence of cognitive bias in raters’ assessments of qualitative statements about language based on the length of the responses (Santhanam et al., 2020; Park et al., 2024). Therefore, we introduced a control by suggesting shorter messages in the system prompt and the max_tokens parameter of the generation pipeline, resulting in the length-controlled baseline with mean length of 75 characters. Interestingly, only some of the emotion scores in the human evaluation dimensions correlate with the generated outputs’ length. See Table 7 for detailed analysis of effect sizes.
@@ 5 Conclusion
We have successfully aligned large language models (LLMs) for empathetic response generation using preference datasets and Direct Preference Optimization (DPO). Our approach demonstrated that preference data can be effectively constructed using Plutchik’s polar emotion model. Through extensive experimentation, we highlighted how varying hyperparameters can mitigate the specialization effect of training and influence empathy metrics. Our findings, validated through the EmpatheticDialogues dataset, diff-Epitome, BERTscore, and diversity and emotion valence automated metrics, as well as human evaluations, provide a solid foundation for further advancements in the preference-based alignment of LLMs for empathetic response generation.
@@
While our method certainly improves the congruence of responses with the ground truth, we believe that training solely on the EmpatheticDialogues dataset alone is not going to fulfill the ultimate goal of aligning LLMs for ERG: generating more empathetic responses in all domains. However, we believe that we have shown the potential of polar emotion models to guide preference dataset generation, on which future work can build.
In human evaluation, we have encountered the well-documented human preference for longer responses Santhanam et al. (2020); Park et al. (2024), which should not be disregarded. Future work should, therefore, focus on combining both insights from our work: collecting or generating long responses while guiding the preference dataset construction using the polar emotion models.
@@ Appendix A Training Details
Code
Models
Training Settings
We trained the models using the HuggingFace Transformers, TRL, and PEFT libraries. In the SFT step, we applied standard sequence-to-sequence training with cross-entropy loss on tokens from the user prompts (not on the tokens from "replies"). In both the SFT and DPO training steps, we optimized the model with AdamW optimizer and batch size of 64. We used learning rate of 1e-5 for the SFT and 1e-6 for DPO without warmup and with a cosine lr decay. The models were trained in bf16 precision.
Hardware
To train our models, we used three NVIDIA A100 80GB GPUs and four A100 40GB GPUs. We trained LoRA adapters with ranks ranging from 16 to 2048 for models with 7B parameters. The total training wall time, including preliminary experiments, was 33 days.
EmpatheticDialogues
We use the version from hf.co/datasets/empathetic_dialogues filtering out $<1\%$ malformed dialogues by keywords. For details we refer the reader to the code repository functions ed_load.{load_preprocess_ed, get_ed_for_generation, get_ed_for_dpo}.
DPO Preference Dataset
@@
Appendix B Open LLM Leaderboard Tasks
Here we describe the tasks what make up the Open LLM leaderboard based on the documentation Beeching et al. (2023b).
Big Bench Hard (BBH)
Suzgun et al. (2023) is a subset of 23 challenging tasks from the BigBench dataset to evaluate language models. The tasks use objective metrics, are highly difficult, and have sufficient sample sizes for statistical significance. They include multistep arithmetic, algorithmic reasoning (e.g., boolean expressions, SVG shapes), language understanding (e.g., sarcasm detection, name disambiguation), and world knowledge. BBH performance correlates well with human preferences, providing valuable insights into model capabilities.
MATH
Hendrycks et al. (2021) is a compilation of high-school level competition problems gathered from several sources, formatted consistently. Generations must fit a very specific output format. A subset of level 5 MATH questions only is included in the benchmark.
Graduate-Level Google-Proof Q&A Benchmark (GPQA)
Rein et al. (2023) is a highly challenging knowledge dataset with questions crafted by PhD-level domain experts in fields like biology, physics, and chemistry. These questions are designed to be difficult for laypersons but relatively easy for experts. The dataset has undergone multiple rounds of validation to ensure both difficulty and factual accuracy. Access to GPQA is restricted through gating mechanisms to minimize the risk of data contamination.
Multistep Soft Reasoning (MuSR)
Sprague et al. (2023) is a dataset consisting of algorithmically generated complex problems, each around 1,000 words in length. The problems include murder mysteries, object placement questions, and team allocation optimizations. Solving these problems requires models to integrate reasoning with long-range context parsing. Few models achieve better than random performance on this dataset.
Massive Multitask Language Understanding - Professional (MMLU-Pro)
Wang et al. (2024b) is a refined version of the MMLU Hendrycks et al. (2020) dataset, which has been a standard for multiple-choice knowledge assessment. Recent research identified issues with the original MMLU, such as noisy data (some unanswerable questions) and decreasing difficulty due to advances in model capabilities and increased data contamination. MMLU-Pro addresses these issues by presenting models with 10 choices instead of 4, requiring reasoning on more questions, and undergoing expert review to reduce noise. As a result, MMLU-Pro is of higher quality and currently more challenging than the original.
IFEval
Zhou et al. (2023) is a dataset designed to test a model’s ability to follow explicit instructions, such as "include keyword x" or "use format y". The focus is on the model’s adherence to formatting instructions rather than the content generated, allowing for the use of strict and rigorous metrics.
Appendix C Automatic Empathy Evaluation: Definitions and Implementation
In this section, we detail the automatic empathy evaluation methods, including a brief definition, method of evaluation, and some discussion on the impact of the metric on the intepretation of the results. We used automatic empathy evaluation methods detailed in Lee et al. (2024), with the exception of the PAIR metric (Min et al., 2022), which we were not able to reliably reproduce.
Specificity
In the patient-practitioner context, specificity is defined as the degree to which the practitioner comments on the generalities versus specific emotions. Lee et al. (2024)’s approach uses normalized inverse document frequency (NIDF) to determine the specificity of a dialogue, as it is a normalized ratio of how rare a word is in a given text.
Word Choice
When evaluating empathy, language which is both rich and consistent with the conversant’s discourse is perceived to be more empathetic. The deviation within the dimensional aspects of emotion (intensity, valence, and arousal) between interactions is used as a metric to gauge the emotional consistentcy in word choice. The NRC Emotion Intensity and NRC-VAD Lexicons were used as the reference.
Here we present the distributions of valence, arousal, and intensity across our models from §3. A distribution for each metric is shown in Figure 9. We provide this distribution to highlight that there are two possible intepretations of the IVA deviation metric: the absolute distance or the signed difference. Lee et al. (2024) adopts the former approach.
Refer to caption Refer to caption Refer to caption
Figure 9: Valence, intensity, and arousal distributions for the two models, two baselines, and human responses.
Furthemore, we present the absolute values of the IVA metrics from §4 in Figure 10. The signed difference approach provides more detailed insight into emotion patterns in word choice, in addition to those presented in the main body of the paper.
Refer to caption Refer to caption Refer to caption
Figure 10: Results of word choice metrics: intensity, valence, and arousal. Distances between the prompt and generated response utterance scores for emotional intensity, valence, and arousal, i.e., $|\text{score}(\text{prompt})-\text{score}(\text{response})|$.
Diversity
Instrumental to the generalizability of empathy evaluation, diversity aims to examine a mode’s capability to generate new responses based on a single utterance, so as to ensure that the model is generating distinct responses rather than a set of templates. Lee et al. (2024) develops a method to examine diversity based on response-tries, i.e. a folded token-bu-token trie representation of the responses provided by the model on repeated prompts. These response tries are constructed by “folding” over a response whenever a unique token is observed. Once prompted for a test set, the properties of the constructed tries, such as the number of templates (i.e. the number of distinct response structures) and compression ratio (the ratio of the size of the response tries before and after folding).
Appendix D Design of Human Evaluation Survey
Human evaluation of EmPO was done on the Prolific platform (Prolific, 2024), by annotators screened for their education level and experience in Psychology and Social Work.
Cohen’s d
Emotion Situational Contextual Conversational
Output source Greater score than Understanding Appropriateness Naturalness Engagingness
baseline: Human targets $0.58$* $0.88$* $0.48$* $0.47$*
baseline-len-control $0.39$* 0.420.42$\dagger$0.42 † 0.30.3$\ddagger$0.3 ‡ $0.45$*
SFT $0.5$* $0.72$* 0.420.42$\dagger$0.42 † $0.53$*
SFT + DPO 0.380.38$\dagger$0.38 † $0.7$* $0.17$ 0.30.3$\dagger$0.3 †
SFT + DPO Human targets $0.18$ $0.14$ 0.320.32$\ddagger$0.32 ‡ 0.180.18$\ddagger$0.18 ‡
baseline-len-control $-0.04$ $0.25$* $-0.16$ $-0.17$
SFT $0.2$ 00 0.250.25$\ddagger$0.25 ‡ 0.220.$2$
baseline-len-control Human targets $0.14$ 0.410.41$\dagger$0.41 † $0.15$ 00
SFT $0.06$ 0.260.26$\ddagger$0.26 ‡ $0.08$ $0.04$
SFT Human targets $0.19$ $0.01$ $0.01$ $-0.01$
Table 7: We used a t-test to test a one-sided hypothesis that the annotators rate some model outputs (left) higher than others (right), * with $p<0.00001$, $\dagger$ with $p<0.001$, and $\ddagger$ with $p<0.05$. After the Bonferronni correction most of the results are not significant on $\alpha=0.05$. But the difference between the corpus-given targets and the baseline is suggesting a bias in the survey design, sub-optimal empathy decomposition coupled with non-ideal dimension definitions, or an issue with the EmpatheticDialogues dataset itself as suggested by Debnath and Conlan (2023).
Existing human evaluations are often done on a discrete Likert scale, along with A/B testing of responses (see Table 1). Current benchmark evaluations use a three-part test of relevance, fluency, and empathy/sympathy Rashkin et al. (2019). As this neither expounds on the multifaceted nature of empathy, nor does it provide a tractable comparison for models, extant literature has adopted several other facets considered important in an empathetic conversational agent (i.e. informativeness, safety, etc.). For our work, we deemed A/B testing unsuitable, as the models generate semantically and situationally similar responses, inducing the cognitive bias towards preferring sentences of longer lengths as opposed to the more “empathetic” ones. We also preferred a continuous 0-100 rating scale in order to maintain tractability Ji et al. (2022).
We use a four-dimensional framework to evaluate empathy in the language, i.e. the design was chosen to best reflect the perceived empathy of the model based on a single conversation. In Table 8, we detail the parameters, displayed statements, and an examination of its choice as a parameter.
As emotions, empathy, and conversation are dynamic, a static evaluation is the aggregate perception of the annotators’ examination of a single instance of a conversation. Therefore, questions about the nature of the model were avoided in favour of response about the nature of the response, as their rating of empathetic measures have to be explicitly based on the language generated by the model and any cues therein.
The human evaluation included two pilot studies. Both pilot studies included 90 participants, recruited using Prolific under the same constraints as the final human evaluations. Finally, the ground truth responses were evaluated as a control.
When extracting examples for the evaluation we first grouped them, then randomly sampled four from each. The conversations were shown to annotators in the same order, but we block-randomized the samples, so that some annotators evaluated outputs of more than one model, in batches of 15. For DPO, base models and human targets, the randomization scheme is as follows: block the outputs, so that 5 examples in a row are from the same model resulting in $3!=6$ orderings. Adding $3$ batches that evaluate samples of one model, making each sample annotated three times. For the length-controlled baseline, we skipped the block-randomization.
Figure 11: The evaluation page for annotation. Annotators were additionally shown an introductory page with in-depth instructions and a solved example with the same structure.
Parameter Statement Explanation
Emotion Understanding The listener’s response addresses the speaker’s emotional state, specifically Emotion understanding is essential for empathetic engagement, but the agent’s comprehension of the user’s emotions can only be assessed if the response explicitly addresses them.
Situational Appropriateness The listener’s response is appropriate to the speaker’s situation. We examine situational appropriateness because it allows for the consideration of context in empathetic responses, unlike a single prompt. Additionally, the ED corpus uses a situational model where conversations are linked to specific emotions and situations. Thus, we assess how well the annotator perceives the agent’s understanding of the situation.
Contextual Naturalness The listener’s response is completely natural in the context of this conversation. Human conversation reflects understanding of emotional and situational context through the nature and tone of responses. This question aims to assess how closely the agent’s response mirrors a human response in a similar context.
Conversational Engagingness I would like to continue the conversation with this listener. This question focuses on the user’s willingness to continue the conversation with the agent, rather than the conversation itself. Given empathy’s role in fostering user-agent bonding, we assess whether annotators would like to continue. However, unwillingness may stem from factors beyond the agent’s engagingness.
Consistency & Fluency The listener’s responses were both understandable and consistent with the rest of the dialogue (yes/no). Relevance and fluency are a staple part of dialogue evaluation. However, both with the advancement of LLMs in general, and the similar base model for our examination in particular, model responses were consistently found to be both consistent and fluent, through both pilots.
Table 8: The parameters of human evaluation with their statement (as provided to the annotators when given instructions), and a brief explanation of why the parameter was selected and the rationale for framing the evaluation statements.